### v3

# Uncertainty-Aware Probabilistic Graph Neural Networks for Road-Level Traffic Accident Prediction

Abstract
Traffic accidents present substantial challenges to human safety and socio-economic development in urban areas. Developing a reliable and responsible traffic accident prediction model is crucial to addressing growing public safety concerns and enhancing the safety of urban mobility systems. Traditional methods face limitations at fine spatiotemporal scales due to the sporadic nature of high-risk accidents and the predominance of non-accident characteristics. Furthermore, while most current models show promising occurrence prediction, they overlook the uncertainties arising from the inherent nature of accidents, and then fail to adequately map the hierarchical ranking of accident risk values for more precise insights. To address these issues, we introduce the Spatiotemporal Zero-Inflated Tweedie Graph Neural Networks (STZITD-GNN), the first uncertainty-aware probabilistic graph deep learning model in road-level traffic accident prediction for multi-steps. This model integrates the interpretability of the statistical Tweedie family model and the expressive power of graph neural networks. Its decoder innovatively employs a compound Tweedie model—a Poisson distribution to model the frequency of accident occurrences and a Gamma distribution to assess injury severity, supplemented by a zero-inflated component to effectively identify exessive non-incident instances. Empirical tests using real-world traffic data from London, UK, demonstrate that the STZITD-GNN surpasses other baseline models across multiple benchmarks and metrics, including accident risk value prediction, uncertainty minimisation, non-accident road identification and accident occurrence accuracy. Our study demonstrates that STZTID-GNN can effectively inform targeted road monitoring, thereby improving urban road safety strategies.
keywords:
Spatiotemporal Sparse Data Mining , Traffic Accident Prediction , Uncertainty Quantification , Zero-Inflated Tweedie Distribution
[…]
[inst1]organization=SpaceTimeLab, University College London (UCL), state=London, country=UK
[…]
[inst2]organization=School of Computer Science, Peking University (PKU), state=Beijing, country=China
[…]
1 Introduction
Traffic accidents persist as a major barrier to sustainable urban development (Dai et al.,, 2018; Tu et al.,, 2020; Barahimi et al.,, 2021; Rodrigue,, 2020). The recent World Health Organisation (WHO) report reveals that road traffic accidents claimed approximately 1.19 million lives worldwide in 2021, making them the leading cause of mortality and disability and imposing an economic burden of up to US\$1.8 trillion, approximately 10–12% of the global Gross Domestic Product (GDP)(Organization,, 2023).
To support reductions in traffic accidents, there is a growing emphasis on the development of models aimed at predicting the location, severity, and causes of accidents. The early approaches relied mainly on linear regression models (Mountain et al.,, 1996; Greibe,, 2003), models that incorporate random parameters (El-Basyouny and Sayed,, 2009; Dinu and Veeraragavan,, 2011), and geographically weighted regression (GWR) models (Li et al.,, 2013; Xu and Huang,, 2015) to predict regional traffic accidents. Recent research has turned to deep learning-based methods, notably graph neural networks (GNNs), which are able to handle large datasets and capture complex spatial and temporal representations of urban road networks. Examples include time-varying GNN (Zhou et al., 2020b, ), multiview fusion GNN(Trirat et al.,, 2023) and subgraph-based GNN(Zhao et al.,, 2023). Despite substantial advancements in model development, accurately predicting road accidents at a fine spatiotemporal granularity road presents several challenges:
  • Spatially Imbalanced Accident Occurrence:The prediction of road accidents is challenging due to the uneven spatial distribution of accidents. Firstly, zero-inflation is inherent to accident data, where areas with no accidents are much more common than those with accidents. This spatially skewed distribution causes models to be biased toward predicting zero accidents, complicating the effective training of predictive algorithms(Wu et al.,, 2022). Secondly, where accidents do occur, they are predominantly minor, leading to an abundance of low-risk scores across the spatial grid. This creates a misleading representation that underestimates the potential severity of less frequent, more severe accidents(Shirazi and Lord,, 2019; Saha et al.,, 2020). The combination of these factors—widespread zero-accident areas and the dominance of minor accidents in accident-prone zones—results in a challenging scenario for models that need to understand and predict the full spectrum of accident severities across different areas (Wang et al.,, 2023).
  • Temporal Constraints with Uncertainty: Building on the spatial challenges, the temporal dynamics of traffic accidents introduce additional complexity. Particularly, predictive models for traffic accidents face the challenge of accurately forecasting not just immediate but also future conditions, necessitating reliable multi-step predictions (Soltani and Qadikolaei,, 2024). The nature of traffic accidents, characterized by their non-linearity and the variable time intervals between events, adds complexity to this task. As the prediction horizon extends, uncertainty increases, particularly because future traffic conditions and potential accident triggers can change substantially over time(Qian et al.,, 2022). Most current deep learning methods do not fully address these challenges. They tend to overlook the complexities of aleatoric uncertainty, which arises from the inherent randomness in the data. These models often rely on assumptions of uniform variance across all data points, producing point estimates from average outcomes(Wang et al.,, 2023; Zhuang et al.,, 2022; Jiang et al.,, 2023). This approach fails to capture the temporal variability and the unpredictable nature of accidents, resulting in predictions that do not adequately reflect the true range of future possibilities.
To address the intricate challenges posed by the spatial and temporal distribution of accident data, we introduce the Spatio-temporal Zero-Inflated Tweedie Graph Neural Network (STZITD-GNN). This model combines spatio-temporal graph machine learning with statistical methods, enabling enhanced uncertainty quantification over multi-step prediction intervals. It is specifically designed to handle the probabilistic assessment of non-recurrent, road-level traffic accidents, providing a robust solution to predict spatially and temporally varying accident data.
The proposed framework integrates a cohesive encoder that combines a Gated Recurrent Unit (GRU) for capturing temporal dynamics and a Graph Attention Network (GAT) for spatial relationships, ensuring a synchronised capture of spatio-temporal dependencies across individual roads and time steps. The decoder leverages a Tweedie (TD)-based distribution, a flexible compound Poisson-gamma model that simultaneously models both exact zero occurrences and continuous positive values. The choice of the TD distribution is motivated by its established efficacy in fields challenged by nonnormal distributions, such as accident frequency estimation (Wang et al.,, 2019), insurance claims (Smyth and Jørgensen,, 2002), actuarial studies (Shi,, 2016), and meteorological precipitation (Dunn,, 2004). This approach allows the model to output the TD distribution parameters, explicitly determining the upper and lower bounds for the prediction intervals of each sample, thus providing a probabilistic representation of each road’s accident risk scores.
Given the inherent extreme imbalance that traditional TD models struggle to adequately capture (Khosravi et al.,, 2010; Abe and Yadohisa,, 2017), we have enhanced our model by incorporating an additional parameter, $\pi$, to estimate the probability mass of zero values. This adjustment leads to the creation of the Zero-Inflated Tweedie (ZITD) distribution on the decoder, which offers a more comprehensive and responsible depiction of real accident scenarios. This refinement ensures that our model not only captures the complexities of traffic accident data, but also enhances the accuracy and reliability of its predictions. Specifically, our main contributions can be summarised as follows.
  • To address the limitations of ST-GNN in modelling complex spatial and temporal relationships, our framework introduces an end-to-end spatio-temporal encoder. This encoder is specifically designed to simultaneously model pairwise relationships and integrate them into a cohesive embedding. This approach enhances the ability of our model to capture and synthesize the intricate dynamics between spatial and temporal data points.
  • An innovative ZITD distribution is employed within our model to effectively address the uncertainties associated with zero-inflation and the long-tailed characteristics prevalent in road accident data. This approach enhances the model’s capability to accurately represent the complex distribution patterns observed in traffic accident data.
  • To the best of our knowledge, the STZITD-GNN is the first work to use probabilistic framework for granular, multistep road-level accident forecasting. The use of Likelihood-based Loss minimization rather than the tradional regression ones enhances its uncertainty-aware predictive accuracy and reliability.
  • Extensive testing with real-world data confirms the superior performance of our framework compared to existing state-of-the-art approaches. Notably, it achieves a reduction in regression error of up to 34.60% in point estimation metrics and shows a significant improvement, increasing at least by 55.07% in uncertainty-based metrics.
This manuscript is structured as follows. In Section 2, we review spatiotemporal modelling methodologies in predicting traffic risk. Section 3 provides a detailed description of the proposed model. Section 4 compares our empirical results with state-of-the-art benchmark models, further enhancing our assertions with illustrative visualisations. In Section 5, we conclude the article by highlighting the advantages of our model and suggesting avenues for future improvements.
2 Literature Review
Predicting traffic accidents is essential for urban planning and mobility management. Initially, research in this area predominantly utilised traditional statistical methods and classical machine learning techniques, such as regression models (Chang,, 2005; Lord et al.,, 2005; Caliendo et al.,, 2007; Bergel-Hayat et al.,, 2013), Bayesian networks (Martin et al.,, 2009; Mujalli and De Oña,, 2011; Hossain and Muromachi,, 2012), tree-based models (Chong et al.,, 2004; Chang and Chen,, 2005; Wang et al.,, 2010; Lin et al.,, 2015), and the k-nearest-neighbour method (Sayed and Abdelwahab,, 1998; Tang and Gao,, 2005; Lv et al.,, 2009). These early studies were generally confined to small geographical areas and demonstrated limited effectiveness in capturing the nonlinear and dynamic spatio-temporal patterns of the contributing factors (Chang and Chen,, 2005; Anderson,, 2009; Zhang et al.,, 2014). In addition, these methods often analyse accident data in isolation, neglecting critical interdependencies between different locations (Wang et al.,, 2013; Wang et al., 2021a, ). As a result, their broader applicability to city-wide analysis with large data size was constrained, highlighting the need for more advanced models capable of addressing these limitations.
Recent advances in deep learning techniques have shown promise in urban traffic accident prediction by incorporating spatio-temporal characteristics of the data. Convolutional neural networks (CNNs), recurrent neural networks (RNNs), and graph-based networks, such as graph-convolution networks (GCNs), have demonstrated their effectiveness in capturing intricate spatial patterns, capturing temporal dependencies, and modelling complex interactions (Wang and Cao,, 2021; Liu et al.,, 2020). As summarised in Table 1, Chen et al., (2016) were among the first to explore city level accident prediction employing a stack denoise autoencoder. They integrated human mobility GPS data and historical incident point data at the grid level to map the real-time accident situation in Tokyo. However, while their approach attempted to capture a larger spatial area, it did not consider urban geo-semantic information for precise and long-term accident prediction. Later, Chen et al., (2018) introduced a Stack Denoise Convolutional Autoencoder that incorporated spatial dependencies using stacked CNNs. Nonetheless, both of these studies overlooked the influence of temporal factors. Although RNNs are capable of handling temporal information, they are more suitable for short-term temporal learning (Ren et al.,, 2018). To better address the temporal aspect, Ren et al., (2018) used Long-Short-Term Memory (LSTM) to consider influential temporal factors between several locations. Their work represented the first attempt to incorporate temporal factors into deep learning models for the prediction of city-wide traffic accidents. Similarly, Moosavi et al., (2019) employed LSTM networks for country-wide accident prediction, demonstrating their efficacy in predicting a binary outcome. Building on the LSTM framework, Yuan et al., (2018) proposed a more advanced model known as the Convolutional Long Short-Term Memory (Hetero-ConvLSTM) neural network based on spatial heterogeneous data. This model aimed to jointly address spatial heterogeneity and capture temporal features in state-wide accident prediction. However, their approach relied on pre-defined moving windows based on the spatial features of Iowa, United States, limiting its generalisability to other cities. Bao et al., (2019) integrated CNNs, LSTM as well as ConvLSTM into a spatiotemporal convolutional long short-term memory network (STCL-Net). Their work presented a more practical and feasible approach for the prediction of city-wide traffic accidents. By combining the strengths of these models, STCL-Net effectively captured spatiotemporal dependencies in accident data and influential factors, leading to improved prediction performance.
To capture local and global dynamics with hierarchical spatial information, Zhu et al., (2019) proposed the Deep Spatio-temporal Attention Learning Framework, which incorporates a spatio-temporal attention mechanism to effectively capture the dynamic impact of traffic accidents at different spatial levels. One notable aspect of their work was the emphasis on using real traffic administrative areas instead of manually divided grids. However, this approach does not incorporate spatial relationships, which limits its capacity to model the spatial heterogeneities that characterise traffic accidents. In a similar vein, Huang et al., (2019) introduced the Deep Dynamic Fusion Network (DFN) framework, which aggregates heterogeneous external factors across both spatial and temporal dimensions. Using a temporal aggregation layer, the DFN framework is designed to automatically capture external influences from latent temporal dimensions.
Although the mentioned studies effectively captured temporal and spatial patterns in traffic accident data, yielding promising prediction outcomes, their lack of consideration of latent spatial correlations within urban structures can disrupt regional patterns in practical urban science and lead to prediction inaccuracies. Therefore, Graph-based networks have become a key method for city-wide traffic accident prediction by leveraging spatial graph features from actual urban networks (Xue et al.,, 2022; Cheng et al.,, 2022; Li et al., 2022a, ; Li et al., 2022b, ).
Zhou et al., 2020b developed the RiskOracle framework by constructing an urban graph of regions and integrating the Differential Time-varying Graph Neural Network (DTGN) to identify dynamic correlations between historical traffic conditions and risks, enabling simultaneous predictions of traffic flow and accidents. To tackle zero-inflation in accident data, they introduced a data enhancement strategy based on a priori knowledge (PKDE), defining subregional accident labels for training. This approach adjusts zero accident values to a range between 0 and 1, maintaining the hierarchy of actual accident levels. Expanding on this foundation, Zhou et al., 2022b refined the DTGN framework by embedding an LSTM-based hierarchical sequence learning architecture, accommodating multiscale historical spatio-temporal data for nuanced long-term and short-term predictions. Wang et al., 2021a and Wang et al., 2021b both focus on the integration of geographical and semantic relationships through different graph structures. While Wang et al., 2021a used accident, road, and POI graphs to capture regional semantics and introduced modules for geographical and semantic spatio-temporal correlations, Wang et al., 2021b ) combined attention-based LSTM, CNN, and GCN for multi-view predictions of traffic accidents at various scales. Although both approaches tackled the zero-inflation issue with promising results, they share limitations in addressing dynamic geographical semantics and alignment issues in graph structures. In contrast, Trirat et al., (2023) proposed a multi-view graph neural network incorporating dynamic and static similarity information, offering a more dynamic approach to understanding traffic accidents. Using the Huber loss function, a parameter-based regression loss, their model aimed at robust regression that effectively accommodates zero inflation, showcasing an advanced strategy for precise prediction of urban traffic accidents.
Recent attempts to apply graph-based deep learning for road-level accident prediction have shown promise, but also highlight areas for improvement. Yu et al., (2021) developed a Deep Spatiotemporal Graph Convolutional Network with a unique three-layer network, treating the road graph, spatio-temporal data, and embeddings separately. They tackled zero inflation through undersampling, balancing the dataset between risky and non-risky roads, and framing accident prediction as a binary classification. While this method effectively predicts accidents at the road level, the binary outcome overlooks the varying degrees of accident risk across roads. Similarly, Wu et al., (2022) designed a fully connected dynamic network to predict road-level accidents by focussing on dynamic spatio-temporal correlations. To mitigate zero inflation, they introduced a cost-sensitive learning module that enhances positive sample classification accuracy by considering the imbalance between zero and non-zero accident values. However, this work only considered limited road segments in New York City with binary accident occurrence outputs, highlighting the need for further development to capture the full spectrum of accident severity and provide a more detailed understanding of road-level accident scenarios.
In summary, despite significant progress in urban traffic accident prediction yielding promising results, existing studies have yet to fully address uncertainty quantification in their predictions and provide multi-step accident assessments at the detailed road level.
Table 1: Summary of Deep Learning Methods in Traffic Accident Prediction
Reference Methodology Spatial Unit Zero-Inflation Issue Prediction Steps Prediction Outcome
Chen et al., (2016) Stack Denoise Autoencoder Regions N/A Single Accident Risk Score
Ren et al., (2018) Long Short-term Memory network Regions N/A Multiple Accident Risk Score
Chen et al., (2018) Stack Denoise Convolutional Autoencoder Regions N/A Single Accident Risk Score
Yuan et al., (2018) Convolutional Long Short-Term Memory Regions N/A Multiple Accident Risk Score
Moosavi et al., (2019) Long Short-term Memory Network Regions N/A Single Accident Occurrence
Zhu et al., (2019) Deep Spatial-Temporal Attention Learning Framework Regions N/A Multiple Accident Risk Score
Huang et al., (2019) Dynamic Fusion Network Regions N/A Single Accident Occurrence
Bao et al., (2019) Spatiotemporal Convolutional LSTM Regions N/A Single Accident Risk Score
Zhou et al., 2020b DTGN with PKDE Regions Yes Single Accident Risk Score
Wang et al., 2021a GSNet with Weighted Loss Regions Yes Single Accident Risk Score
Wang et al., 2021b MVMT-ST Networks Regions Yes Single Accident Risk Score
Yu et al., (2021) Deep ST-GCN with Undersampling Roads Yes Single Accident Occurrence
Wu et al., (2022) MADGCN with Cost-sensitive Loss Roads Yes Multiple Accident Occurrence
Zhou et al., 2022b Enhanced DTGN Regions Yes Multiple Accident Risk Score
Trirat et al., (2023) MV-GNNs with Huber Loss Regions Yes Multiple Accident Risk Score
3 Methodology
In this section, we first introduce the concepts and define the framework for predicting road-level traffic accidents in Section 3.1. We then illustrate the statistical innovations from the TD distribution to ZITD in the context of uncertainty quantification in Section 3.2. Subsequently, we introduce our new STZITD-GNN model, a spatio-temporal probabilistic modelling framework that combines the strengths of GRU and GAT in Section 3.3. The section ends with an in-depth exploration of the loss function employed within our model in Section 3.4. The major notations used in this section are listed in Table 2.
Table 2: Main notations in this paper
Notation Description
$Y\leavevmode\nobreak\ /\leavevmode\nobreak\ X$ Road-level traffic accident risk score  /  Road-level spatiotemporal traffic feature
$V\leavevmode\nobreak\ /\leavevmode\nobreak\ E\leavevmode\nobreak\ /\leavevmode\nobreak\ A$ Road set  /  Edge set  /  Road adjacency matrix
$v_{i}\leavevmode\nobreak\ /\leavevmode\nobreak\ (v_{i},v_{j})$ The $i^{th}$ road  /  Edge between the $i^{th}$ road and $j^{th}$.
$N\leavevmode\nobreak\ /\leavevmode\nobreak\ T$ Numbers of roads  /  Time length
$C_{it}$ Number of accidents on road $v_{i}$ at time slot $t$
$l_{it}^{j}$ The $j^{th}$ cardinality values of accident severities in the time slot $t$ for road $v_{i}$
$d\leavevmode\nobreak\ /\leavevmode\nobreak\ F\leavevmode\nobreak\ /\leavevmode\nobreak\ F^{\prime}$ Feature dimension/Temporal feature dimension/Spatio-Temporal feature dimension
$\mathcal{ST}$ Spatiotemporal deep learning Model
$f_{\text{TD}}(\cdot)$ Probabilistic density function of Tweedie distribution
$\rho/\leavevmode\nobreak\ \phi/\leavevmode\nobreak\ \mu/\leavevmode\nobreak\ \pi$ Index parameter/ Dispersion parameter/ Mean parameter/ Zero-inflated parameter
$\theta\leavevmode\nobreak\ /\leavevmode\nobreak\ \lambda\leavevmode\nobreak\ /\leavevmode\nobreak\ \alpha,\gamma$ Natural parameter  /  Poisson mean parameter  /  Gamma parameters
$\mathcal{Z}_{T}\leavevmode\nobreak\ /\leavevmode\nobreak\ \mathcal{Z}$ Temporal embedding  /  Spatiotemporal embedding
$\text{ReLU}(\cdot)/\leavevmode\nobreak\ \sigma(\cdot)/\leavevmode\nobreak\ \text{LeakyReLU}(\cdot)/\leavevmode\nobreak\ [\cdot||\cdot]$ ReLU activation / Sigmoid activation / LeakyReLU activation / Concatenation
$W_{r},W_{u},W_{c},W_{a},b_{r},b_{u},b_{c},\vec{a}$ Learnable weight matrics of GRU and GAT
$W_{\pi},W_{\mu},W_{\phi},W_{\rho},b_{\pi},b_{\mu},b_{\phi},b_{\rho}$ Learnable weight matrics of the four Parameter Encoders
$h_{t}\leavevmode\nobreak\ /\leavevmode\nobreak\ r_{t}\leavevmode\nobreak\ /\leavevmode\nobreak\ u_{t}$ Hidden feature  /  Reset gate  /  Update gate
$\alpha_{i,j}$ Attention value between road $v_{i}$ and $v_{j}$
$M$ Number of attention heads
$\epsilon$ The minimum value
$\eta$ L2 normalisation weight-parameter
3.1 Preliminaries and Problem Definition
This section presents the basic concepts and terminology relevant to the study, followed by a description of the problem.
Definition 1. Road Traffic Accident Risk Score: A traffic accident refers to an accident that involves one or more road users that leads to physical injury, loss of life, or property damage. We denote the set of roads as $V$ of size $N=|V|$, where $v_{i}$ signifies the $i^{th}$ road. The traffic accident risk on the roads is represented as $Y\in\mathbb{R}^{N\times T}$ for a duration of total $T$ periods, and $y_{it}$ denotes the accident risk in the $t^{th}$ time slot for the road $v_{i}$. Thus, $Y_{t}\in\mathbb{N}^{N}$ signifies the accident risk for all roads in the $t^{th}$ time interval, where $y_{it}$ is its constituent element. The calculation of traffic accident risk score $y_{it}$ for the road $v_{i}$ at the time slot $t$ is given by:
$y_{it}=\sum_{j=1}^{3}C_{i,k}^{t}\times l_{j},$ (1)
where $C_{i,k}^{t}$ denotes the total number of $k$ accidents with the severity level $l_{j}$ at time slot $t$ for road $v_{i}$. We allocate the accident point to its closest road. As per prior research (Wang et al., 2021a, ; Trirat et al.,, 2023), $l$ is assigned the values 1, 2, and 3, representing minor injury, serious injury, and fatal accident severities, respectively. $y_{it}$ equals 0 if there are no traffic accidents.
Definition 2. Spatio-temporal Features: The embedding features are represented as $X\in\mathbb{R}^{N\times T\times d}$, where $d$ is the dimension of the features. We denote $x_{it}$ as the spatio-temporal feature of the road $v_{i}$ at the time slot $t^{th}$. Consequently, $X_{t}\in\mathbb{R}^{N\times d}$ signifies the features of all roads at time slot $t$. Our study includes spatial features such as road types, road lengths and widths, road conditions, and census characteristics for each lower layer super output area (LSOA) of roads. Temporal features, which vary daily, encompass weather information such as sunrise and sunset times, humidity, visibility, rainfall, etc., along with the identification as a holiday or a working day.
Definition 3. Road Connection Graph: The road network is represented as a graph $\mathcal{G}=(V,E,A)$. Here, $E$ denotes the set of edges and $A\in\mathbb{R}^{N\times N}$ is the adjacency matrix that describes the connections between roads. An entry $A_{i,j}=1$ indicates the existence of an edge between road $v_{i}$ and $v_{j}$, denoted as $(v_{i},v_{j})\in{E}$, while $A_{i,j}=0$ implies no connection, denoted as $(v_{i},v_{j})\notin{E}$.
Problem Definition: Our model aims to predict the future accident risk score in the next $p$ time windows and the confidence interval of the predicted results per road. It is to use historical records $X_{1:t},Y_{1:t}$ and the graph structure $A$ as input data for training, to predict the probabilistic density function $f(Y_{t+1:t+p})$ of the distribution of $Y_{t+1:t+p}$.
$\mathcal{ST}_{\Theta}([X_{1:t},Y_{1:t}],A)\to f(Y_{t+1:t+p}),$ (2)
where $\Theta$ denotes the learned parameters for the ST-GNN model (also referred to as $\mathcal{ST}$). The multistep predictions are made for a 14-day output, as $p=14$.
Figure 1: The overall framework of STZITD-GNNs. STZITD-GNNs utilise the ST-GNN Encoder $\mathcal{ST}$ (composed of GRU and GAT encoders) to encode the history time window $1:t$ accident risk $Y_{1:t}$ and road features $X_{1:t}$ with the use of the road connection graph in the spatial-temporal embedding of the road $\mathcal{Z}$. After encoding, the four parameter decoders map $\mathcal{Z}$ into the ZITD parameter space and obtain $\pi_{t+1:t+p},\mu_{t+1:t+p},\phi_{t+1:t+p},\rho_{t+1:t+p}$ for the predicted time window $t+1:t+p$, which determine the predicted road traffic accident risk distribution $f(Y_{t+1:t+p})$.
3.2 From TD Distribution to ZITD Distribution
This subsection provides an in-depth exploration of the progression from the TD to the ZITD distribution in the specific context of road accident prediction. Building on the work of Lord and Mannering, (2010) and Smyth and Jørgensen, (2002), we recognise the effectiveness of TD distribution to address overdispersed data, indicating its potential applicability in multi-step road-level traffic accident prediction.
As illustrated in Section 3.1, when we aggregate accident data (incorporating various severities) from different roads or identical roads over different time intervals, the resulting distribution of accident counts tends to exhibit overdispersion. In detail, the accident value applied to both accident counts and the associated severity - is the prevalence of ’excess’ zeros, implying more zero entries than would be predicted under a corresponding Poisson process. Meanwhile, in nonzero accident values, a clearly left-skewed but long-tailed distribution presents a smaller proportion of higher severity accidents. In these scenarios, the variance of the accident counts outweighs the mean accident count, creating a deviation from the Gaussian distribution, as underlined by the work of Halder et al., (2019) and Wang et al., (2023).
3.2.1 TD Distribution
In accordance with Equation 1, we designate $y_{k}$ as the accident risk value of the road $i$ on time slot $t$. As a special case of Exponential Dispersion Models (EDMs), the TD model encompasses a broad family of statistical distributions (Tweedie et al.,, 1984; Jørgensen,, 1987; Jiang et al.,, 2023). A random variable that follows the TD distribution has a probability density function $f_{TD}$ (Jørgensen,, 1987), as follows :
$f_{\text{TD}}(y_{k}|\theta,\phi)\equiv a(y_{k},\phi)\exp{\bigl{[}\frac{y_{k}(\theta)-\kappa(\theta)}{\phi}\bigl{]}},$ (3)
here, $\theta\in\mathbb{R}$ represents the natural parameter, while ϕ+italic-$\phi\in\mathbb{R}^{+}$ stands for the dispersion parameter. The normalizing functions $a(\cdot)$ and $\kappa(\cdot)$ are parameter functions for $\phi$ and $\theta$, respectively, detailed further in the subsequent discussion. For a TD distribution of EDMs, the mean and variance of variable $y_{k}$ are given by:
$E(y_{k})=\mu=\kappa(\theta)^{{}^{\prime}},\quad Var(y_{k})=\phi\kappa(\theta)^{{}^{\prime\prime}},$ (4)
In these equations, $\kappa(\theta)^{{}^{\prime}}$ and $\kappa(\theta)^{{}^{\prime\prime}}$ represent the first and second derivatives of $\kappa(\theta)$, respectively. Meanwhile, the function $\mu\geq 0$ works as the mean parameter. Since the mean $E(y_{k})$ and variance $Var(y_{k})$ are dependent, $\kappa(\theta)^{{}^{\prime\prime}}$ can be directly expressed as $V(\mu)$, thereby providing a direct mean-variance relationship (Mallick et al.,, 2022). Additionally, the TD family encompasses a number of vital distributions for distinct index parameters $\rho$ as demonstrated by $V(\mu)=\mu^{\rho}$. These include the Normal ($\rho=0$), Poisson ($\rho=1$), Gamma ($\rho=2$), Inverse Gaussian ($\rho=3$), and Compound Poisson-Gamma distribution ($21<\rho<$ 2) and could be referred to as A. The Compound Poisson-Gamma distribution is particularly applicable, given its focus on parameterizing zero-inflated and long-tailed data, which aligns with the observed characteristics of real traffic accident data.
The frequency of traffic accidents $C_{k}$ adheres to a Poisson distribution represented by $Pois(\lambda)$ with a mean of $\lambda$, expressed as $C_{k}\sim Pois(\lambda)$. This is natural, as frequency often appears as the main prediction target in the previous literature (Zhou et al., 2020b, ; Trirat et al.,, 2023). Currently, the severity of the accident $l_{k}^{(j)}$ is modelled as independent and identically distributed $i.i.d$ Gamma random variables, $C_{k}\perp l_{k}^{(j)}$ and $\forall j$, denoted as $Gamma(\alpha,\gamma)$. These variables have a mean of $\alpha\gamma$ and a variance of $\alpha\gamma^{2}$, such as $l_{k}^{(j)}\stackrel{{\scriptstyle iid}}{{\sim}}\operatorname{Gamma}(\alpha,\gamma)$. $C_{k}$ and $l_{k}^{(j)}$ together form the model of the occurrence of a single accident. Consequently, the traffic accident risk $y_{k}$ is formulated as a Poisson sum of these i.i.d. Gamma random variables. Considering Equation 1 and considering $21<\rho<$ 2, we can finally reformulate the random variable $y_{k}$ as follows:
$y_{k}=\left\{\begin{array}[]{cc}0&\text{if }C_{k}=0,\\ \sum_{j=1}^{C_{k}}l_{k}^{(j)}=l_{k}^{(1)}+l_{k}^{(2)}+\cdots+l_{k}^{(C_{k})}&\text{if }C_{k}>0.\end{array}\right..$ (5)
In the scenario where $C_{k}=0$, we correspondingly have $y_{k}=0$. The probability density at zero for traffic accident risk, therefore, becomes $\mathbb{P}(y_{k}=0)=\mathbb{P}(C_{k}=0)=\exp(-\lambda)$. In contrast, when $C_{k}>0$, $y_{k}$ is deduced from the sum of $C_{k}$ independent Gamma random variables. For clarity, following the work of Zhou et al., 2022a , we re-parameterize the TD distribution, setting $\theta=\mu^{1-\rho}/(1-\rho)$ and $\kappa(\theta)=\mu^{2-\rho}/(2-\rho)$. This results in the revised TD distribution as:
$f_{\text{TD}}(y_{k}|\theta,\phi)=f_{\text{TD}}(y_{k}|\mu,\phi,\rho)=a(y_{k},\phi,\rho)\exp{\bigl{[}\frac{1}{\phi}({y_{k}\frac{\mu^{1-\rho}}{1-\rho}-\frac{\mu^{2-\rho}}{2-\rho}})\bigl{]}},$ (6)
where
$a(y_{k},\phi,\rho)=\left\{\begin{array}[]{cc}1&\text{if }y_{k}=0,\\ \frac{1}{y_{k}}\sum_{j=1}^{\infty}\frac{y_{k}^{-j\alpha}(\rho-1)^{\alpha j}}{\phi^{j(1-\alpha)}(2-\rho)^{j}j!\Gamma(-j\alpha)}&\text{if }y_{k}>0.\end{array}\right..$ (7)
In this case, $\mu,\phi$ and $\rho$ are the three parameters that govern the probability and expected value of the risk of traffic accidents. Parameters in the Gamma and Poisson distributions ($\lambda,\alpha,\gamma$) can be obtained as (Kurz,, 2017; Halder et al.,, 2019; Mallick et al.,, 2022), :
$\lambda=\frac{1}{\phi}\frac{\mu^{2-\rho}}{2-\rho},\quad\alpha=\frac{{2-\rho}}{\rho-1},\quad\gamma=\phi(\rho-1)\mu^{\rho-1}.$ (8)
3.2.2 ZITD Distribution
The TD model, introduced in Section 3.2.1, is designed to account for the occurrence of zeros by allowing a positive probability of zero outcomes, as demonstrated by formula $\mathbb{P}(y_{k}=0)=\mathbb{P}(C_{k}=0)=\exp(-\lambda)$. However, this model struggles to accurately represent accident data on roads where zero values overwhelmingly dominate, leading to an extremely unbalanced distribution that can misrepresent the actual accident rate (Mallick et al.,, 2022; Zhou et al., 2022a, ).
To address this issue, we introduce a refinement, known as the Zero-Inflated Tweedie (ZITD) model with Four-parametric compound distributions. The ZITD model includes a distinctive feature, a sparsity parameter ($\pi$) that specifically addresses the skewness of the data characterised by an excess of zeros. The parameter $\pi$ signifies the likelihood of zero inflation, effectively distinguishing between zero and nonzero occurrences. In contrast, the complement probability ($1-\pi$) aligns with the traditional TD, facilitating a balanced representation of the accident risk values across the spectrum.
$y_{k}=\left\{\begin{array}[]{cl}0&\text{with probability }\pi\\ Y&\text{with probability }1-\pi,Y\sim f_{TD}(Y|\mu,\phi,\rho)\end{array}\right..$ (9)
Finally, denoting the ZITD model as $y_{k}\sim f_{ZITD}(\pi,\mu,\phi,\rho)$, we formulate the probability density function of this model as follows:
$f_{ZITD}(y_{k}|\pi,\mu,\phi,\rho)=\left\{\begin{array}[]{cc}\pi+(1-\pi)f_{TD}(y_{k}=0|\mu,\phi,\rho)&\text{if }y_{k}=0\\ (1-\pi)f_{TD}(y_{k}|\mu,\phi,\rho)&\text{if }y_{k}>0\end{array}\right..$ (10)
In accordance with this specification, we tackle the problem of extreme zero inflation by applying a linear weighting to the zero value, such that $\mathbb{P}(y_{k}=0)=\pi+(1-\pi)\exp{(-\frac{1}{\phi}\frac{\mu^{2-\rho}}{2-\rho})}$. This probability is significantly higher than the probability of observing a zero count from the TD distribution, intuitively indicating a substantial zero inflation. This modification ensures that the mean value of the ZITD distribution is $\mathbb{E}(y_{k})=(1-\pi)\mu$.
In the ZITD model, each parameter plays a distinct role in representing road-level accident risks. The sparsity parameter ($\pi$) is adjusted for instances without accidents, significantly influencing the dispersion parameter ($\phi$) to provide a weighted interpretation suitable for left-skewed accident data. Higher $\phi$ values reflect the improved capability of the model to represent scenarios with lower risk accurately. Additionally, the index parameter ($\rho$) is essential to capture the long-tail distribution of accident risk values. Values near 2 for $\rho$ indicate the robust performance of the model in these scenarios.
3.3 The framework of Spatiotemporal Zero-Inflated Tweedie Graph Neural Networks (STZITD-GNNs)
In relation to the four parameters $\pi,\mu,\phi,\rho$ of the ZITD probability distributions detailed above, we have designed a probabilistic spatio-temporal graph learning module, named STZITD-GNNs. This module extracts the spatiotemporal correlations $\mathcal{Z}=\mathcal{ST}_{\Theta}(X_{1:t},Y_{1:t},A)$ under the assumption of ZITD for each road accident and predicts future results for the subsequent $p$ time windows through $f_{ZITD}(X_{t+1:t+p}|\mathcal{Z})$, as described in Equation 11.
$f_{ZITD}(Y_{t+1:t+p}|\pi_{t+1:t+p},\mu_{t+1:t+p},\phi_{t+1:t+p},\rho_{t+1:t+p})$ (11)
$=f_{ZITD}(Y_{t+1:t+p}|\mathcal{ST}_{\Theta}(X_{1:t},Y_{1:t},A))=f_{ZITD}(X_{t+1:t+p}|\mathcal{Z}).$
As illustrated in Figure 1, instead of performing two independent estimations, our methodology starts with the deployment of a GRU acting as a temporal encoder. Subsequently, a GAT is implemented for the purpose of spatial encoding. The combined output from the GRU and GAT is utilised to parameterize the ZITD distribution. The proposed approach is shown in Equation 12. The architecture’s encoding component applies a spatio-temporal embedding scheme complemented with an extra sparsity parameter and distribution parameters. The decoding component is responsible for probabilistic estimations of future multi-step accident risk values, which can be construed as the set of parameters of the outcome distribution. In contrast to the principles of variational autoencoders, which restrict latent embedding with a specific Gaussian distribution for the representation of latent variables (Kipf and Welling, 2016b, ; Hamilton,, 2020), we echo the statistical domain by allowing integration of deep learning methodologies into parameter learning.
$\pi_{t+1:t+p},\mu_{t+1:t+p},\phi_{t+1:t+p},\rho_{t+1:t+p}=\mathcal{ST}_{\Theta}(X_{1:t},Y_{1:t},A).$ (12)
3.3.1 Parameter Decoders
The detailed functions of the temporal encoder, GRU and the spatial encoder, GAT are listed in the B and C. In this section, we highlight the parameter decoders in detail. To transmute the acquired spatio-temporal embeddings into ZITD parameter values, we have devised four parameter decoders to operate on $\mathcal{Z}$. Consequently, four parameters, namely $\pi_{t+1:t+p},\mu_{t+1:t+p},\phi_{t+1:t+p},$ and $\rho_{t+1:t+p}$ can be computed from $\mathcal{Z}$, as detailed below:
$\pi_{t+1:t+p}$ $=\sigma(W_{\pi}\cdot\mathcal{Z}+b_{\pi})$ (13)
$\mu_{t+1:t+p}$ $=\text{ReLU}(W_{\mu}\cdot\mathcal{Z}+b_{\mu})$
$\phi_{t+1:t+p}$ $=\text{ReLU}(W_{\phi}\cdot\mathcal{Z}+b_{\phi})+\epsilon$
$\rho_{t+1:t+p}$ $=\sigma(W_{\rho}\cdot\mathcal{Z}+b_{\rho})+1+\epsilon$
Here, $\pi$ lies within the range of $[0,1]$, $\mu$ falls within the interval $[0,+\infty)$, $\phi$ exists within $(0,+\infty)$, and $\rho$ spans the range of $(1,2)$. The learnable weight matrices are represented as $W_{\pi},W_{\mu},W_{\phi},W_{\rho},b_{\pi},b_{\mu},b_{\phi},b_{\rho}\in\mathbb{R}^{F^{\prime}\times p}$. $\lim\epsilon\to 0$ represents the smallest value, and $\text{ReLU}(\cdot)$ signifies the ReLU activation function.
3.4 Learning Framework of STZITD-GNNs
3.4.1 ZITD Loss Function
As previously mentioned, our encoding mechanism employs a spatio-temporal embedding architecture complemented by a sparsity parameter. The decoding component involves a probabilistic estimation of future road traffic accidents. To accurately predict these accidents, the overall learning objective can be seen as maximising the log-likelihood function: $\max\log f_{ZITD}(y_{k}|\pi,\mu,\phi,\rho)$. We use the negative likelihood as our loss function to better fit the distribution into the data. Here, $y$ denotes the ground-truth values corresponding to one of the predicted accidents with the parameters $\pi,\mu,\phi,\rho$. The log-likelihood of ZITD is made up of the parts $y=0$ and $y>0$:
For $y>0$, $NLL_{y>0}=-\log f_{ZITD}(y>0|\pi,\mu,\phi,\rho)$:
$\log f_{ZITD}(y>0|\pi,\mu,\phi,\rho)$ $=\log(1-\pi)+\log f_{\text{TD}}(y>0|\mu,\phi,\rho)$ (14)
$=\log(1-\pi)+\frac{1}{\phi}\bigl{(}y\frac{\mu^{1-\rho}}{1-\rho}-\frac{\mu^{2-\rho}}{2-\rho}\bigl{)}+\log a(y>0,\phi,\rho)$
$=\log(1-\pi)+\frac{1}{\phi}\bigl{(}y\frac{\mu^{1-\rho}}{1-\rho}-\frac{\mu^{2-\rho}}{2-\rho}\bigl{)}-\log y+\log\sum_{j=1}^{\infty}\frac{y^{-j\alpha}(\rho-1)^{\alpha j}}{\phi^{j(1-\alpha)}(2-\rho)^{j}j!\Gamma(-j\alpha)}$
$\geq\log(1-\pi)+\frac{1}{\phi}\bigl{(}y\frac{\mu^{1-\rho}}{1-\rho}-\frac{\mu^{2-\rho}}{2-\rho}\bigl{)}-\log(j_{max}\sqrt{-\alpha}y)+j_{max}(\alpha-1).$
In this equation, $j_{max}=\frac{y^{2-\rho}}{(2-\rho)\phi}$ and $\alpha=\frac{2-\rho}{1-\rho}<0$. We optimise the lower bound; hence, $NLL_{y>0}$ can be optimised in this manner.
For $y=0$, $NLL_{y=0}=-\log f_{ZITD}(y=0|\pi,\mu,\phi,\rho)$:
$\log f_{ZITD}(y=0|\pi,\mu,\phi,\rho)$ $=\log(\pi)+\log(1-\pi)f_{\text{TD}}(y=0|\mu,\phi,\rho)$ (15)
$=\log(\pi)+\log(1-\pi)+\frac{1}{\phi}\bigl{(}-\frac{\mu^{2-\rho}}{2-\rho}\bigl{)}$
Here, $\pi,\mu,\phi,\rho$ are chosen and calculated in accordance with the index of $y=0$ or $y>0$. The ultimate negative log-likelihood loss function is given by:
$NLL_{STZITD}=NLL_{y=0}+NLL_{y>0}+\eta{\Theta^{2}}.$ (16)
Model parameter $\Theta$ can be optimized by minimizing the negative log-likelihood loss: $\hat{\Theta}=\text{argmin}NLL_{STZITD}$. $\eta$ stands for the weight parameter for $L_{2}$ Normalisation.
It is worth noting that our model can be extended to other distributions by modifying the probability layer. The main reason is that we learn the parameters that determine the distributions. For example, if we choose the Gaussian distribution, we can parameterize the probability layer using the spatio-temporal embedding of the mean and variance, thereby quantifying the data uncertainty that follows the Gaussian distribution. Using the flexibility of the probability layer, we design a range of benchmark models for comparison to the STZITD-GNNs. Our code is available on Github  [^1] https://github.com/STTDAnonymous/STTD.
3.4.2 Time analysis & Training algorithms
We summarise the STZITD-GNN training procedure in Algorithm 1. Based on that, the time complexity can be analyzed as follows (for the sake of simplicity, we denote the hidden union of embedding $F^{\prime}\approx F$ and $F$ as $F$):
  • For temporal encoders, the time complexity can be formulated as $O\bigl{(}tN(dF+F^{2}+F)\bigl{)}$.
  • For spatial encoders, we assume that the number of road edges is $|E|$. Hence, when the graph attention encoder is implemented, the complexity of computing attention becomes $O(M|E|F)$. Furthermore, when performing $W_{a}$ multiplication, the cost is $O(MNF^{2})$. This leads to a total time complexity of $O(M|E|F+MNF^{2})$.
  • When generating the four parameters, the four one-layer MLPs require a time complexity of $O\bigl{(}F^{2}p\bigl{)}$.
Algorithm 1 The training of STZITD-GNNs.
1:Spatiotemporal road feature $X\in\mathbb{R}^{N\times T\times d}$, road network graph $G=(V,E,A)$, roads traffic accident $Y\in\mathbb{R}^{N\times T}$, model hyperparameters $M,F^{\prime},F$, and total time slot $T$, time window $t$, $p$.
2:Calculate road traffic accident $Y$ via Eq. (1).
3:Initialise parameters $\Theta$ via Xavier Initializer.
4:while STZITD-GNNs is not converged do $\triangleright$ Train
5:    Calculate temporal embedding $\mathcal{Z}_{T}$ via Eq. (22) by road feature $X_{1:t}$; $\triangleright$ Temporal Encode
6:    Calculate spatiotemporal embedding $\mathcal{Z}$ via Eq. (25) by graph $G$ and $\mathcal{Z}_{T}$; $\triangleright$ Spatial Encode
7:    Obtain four parameters $\pi_{t+1:t+p},\mu_{t+1:t+p},\phi_{t+1:t+p},\rho_{t+1:t+p}$ via output layer refer to Eq. (13);
8:    Calculate distribution $f_{ZITD}(y_{t+1:t+p}|\pi_{t+1:t+p},\mu_{t+1:t+p},\phi_{t+1:t+p},\rho_{t+1:t+p})$ from four parameters via Eq. (10);
9:    Minimizing negative log-likelihood loss ${NLL}_{STZITD}$ against $Y_{t+1:t+p}$ via Eq. (16) using Adam Optimizer;
10:end while
11:End optimizing model parameters $\Theta$;
4 Experiments
In this section, we describe the detailed elements of the STZITD-GNNs model by analysing its performance in accident prediction in three boroughs of London, UK, namely Lambeth, Tower Hamlets and Westminster, each characterised by distinct sociodemographic traits, road networks, and traffic accident profiles. The accident data is accessed from the UK STATS19 dataset, maintained by the Department for Transport. It is made up of three interlinked tables that provide information about accidents, their severity, and related vehicle data  [^2] https://www.gov.uk/guidance/road-accident-and-safety-statistics-guidance. We have configured the model to forecast 14 steps ahead, achieving stable temporal horizon results. Our evaluation focusses on two key aspects: first, demonstrating the model’s ability to effectively quantify the uncertainty of road-level traffic accident data, and second, showing its precision in predicting. We emphasise that our model not only forecasts multi-step accident risk in advance but also does so with a higher degree of reliability and precision at the finer granularity compared to several relevant baseline models.
Section 4.1 introduces the datasets used in this study and justifies the choice of data-driven strategies. This is followed by an outline of the experimental setup in Section 4.2. Subsequently, we describe the evaluation metrics in Section 4.3 and the baseline models used for comparison in Section 4.4. The subsequent analysis includes a comparison of the performance of our model against other methods in Section 4.5. It concludes with a sensitivity analysis in Section 4.6
4.1 Traffic Accident Data Description
4.1.1 Statistic Analysis
(a) The map of three boroughs, London
(b) The distribution of accident risk scores frequency in the three boroughs
Figure 2: Traffic accidents are depicted for the case studies.
The three high-accident frequency boroughs of Greater London, UK are illustrated in Figure 2. These boroughs each averaged 3 to 5 daily accidents throughout 2019, as detailed in Table 3. Despite these figures, a significant number of roads in these boroughs report no accidents, leading to a pronounced zero-inflation problem. This situation is compounded by the distribution of accident risk values (Figure 2), which shows a zero-inflated problem along with a pronounced long-tailed effect -where a few high-risk accident occurrences contrast with many low-risk ones. Such data characteristics underscore the suitability of our proposed model for this dataset, effectively capturing both the prevalent zero rates and the diverse spread of accident risk levels.
Table 3: The statistics of Westminster, Lambeth, and Tower Hamlets datasets.
Dataset  Roads/Nodes   Edges   Accident Counts   Zero-Inflated
[…]
  • ${}^{\text{1}}$ https://roadtraffic.dft.gov.uk/downloads
4.1.2 Lower-Dimensional Representation for Traffic Accident
To interpret the representation learnt from the spatiotemporal encoder, we use the Multidimensional Scaling (MDS) algorithm to reduce high-dimensional data to two dimensions while preserving inherent structural information and latent relationship (Wickelmaier,, 2003). This technique facilitates an intuitive visual representation of the data learnt by the deep learning models. For visualisation, we depict nonaccident and low-risk roads with green nodes, while blue nodes represent high-risk roads. The spatiotemporal embeddings, $\mathcal{Z}$, are illustrated in Figure 3.
(a) Lambeth
(b) Tower Hamlets
(c) Westminster
Figure 3: Low accident risk (green color) and high accident risk (blue color) spatio-temporal embedding $\mathcal{Z}$ MDS-visualisation on Lambeth, Tower Hamlets and Westminster with multi-step predict.
This visualisation reveals that roads associated with accidents are naturally clustered yet overlap with some low-risk areas, which introduces aleatoric uncertainties to the representation learning. Thus, traditional models may not be suitable in this case (Zhuang et al.,, 2022). Furthermore, this overlap can be attributed to spatial spillover effects, where adjacent roads exhibit similar embeddings despite their different risk levels. Although low- and high-risk roads with accident clusters are not distinct, high-risk roads within the Westminster data set are more discernible. Both the Lambeth and Tower Hamlets datasets show spatial aggregations that partially reflect the severity of accident risk, indicating that the requirements for an advanced model are to effectively incorporate accident risk severity in its embeddings and to be capable of capturing the complex long-tail distribution of risk within the accident data.
4.2 Experiment Setup
The parameters of STZITD-GNN are optimised with Adam optimiser (Kingma and Ba,, 2014), where the learning rate is set to 0.01, and weight decay is 0.01. In STZITD-GNN, $N_{epoch}=20$, the hidden unit is set to 42, GAT head is set to 3. The baseline parameters are optimised using Adam Kingma and Ba, (2014) with regularisation of $L_{2}$ and a dropout rate of 0.2. The GNNs in STZITD-GNNs and baselines are all two-layered. We also set the hidden units at 42. Similarly to previous work, we also employ the early stopping strategy with patience equal to 10 to avoid overfitting(Zhou et al., 2020a, ). STZITD-GNNs is implemented in Pytorch 1.9.0 with Python 3.8. All experiments were carried out on 1 NVIDIA GeForce RTX 3090, 24 GB.
4.3 Evaluation Metrics
We use eight metrics to evaluate the performance of our STZITD-GNNs from four perspectives, including the predictive regression effect of road accident risk scores, the predictive effect of uncertainty-aware distributional characteristics, the accuracy of zero accident road identification, and the accurate identification of accidents occurring on predicted high-risk roads.
4.3.1 Accident Accuracy Metrics
For the aspect of the point estimate, we compare the prediction of the numerical accident risk scores. For probabilistic distribution outputs like our model, we use the mean value of the distribution as the numerical output. For specific metrics, we use Mean Absolute Error (MAE), Mean Absolute Percentage error (MAPE), and Root Mean Squared Error (RMSE) to evaluate the predicted accident risk scores of each road. Specific definitions are summarized:
$\text{MAE}=\frac{1}{pN}\sum_{j=1}^{p}\sum_{i=1}^{N}\left|y_{ij}-\hat{y}_{ij}\right|,$ (17)
$\text{MAPE}=\frac{1}{pN}\sum_{j=1}^{p}\sum_{i=1}^{N}\left|\frac{y_{ij}-\hat{y}_{ij}}{y_{ij}}\right|,$
$\text{RMSE}=\sqrt{\frac{1}{pN}\sum_{j=1}^{p}\sum_{i=1}^{N}(y_{ij}-\hat{y}_{ij})^{2}},$
where $\hat{y}_{ij}$ and $y_{ij}$ are the predicted (i.e. distribution mean) and ground-truth values of $i^{th}$ road at time slot $j^{th}$ respectively. $p=14$ days, and the time interval $j^{th}$ is the predicted window index. Lower values of these metrics mean better performance with fewer errors.
4.3.2 Uncertainty Quantification Metric
To quantify the uncertainty of the results, we use two typical metrics: the mean prediction interval width (MPIW) and the prediction interval coverage probability (PICP) are based on a confidence interval 5%-95% with a confidence level set at 5% ($\alpha=5\%$). MPIW gauges the average width of the prediction interval, reflecting the distribution’s specific characteristics. Ideally, a smaller MPIW indicates reduced variance and increased certainty, highlighting a more precise prediction. The PICP, on the other hand, is implemented to evaluate whether the ground truth value, denoted $y_{ij}$, lies within the predicted interval. A higher PICP indicates better performance. The goal is to achieve a prediction interval that is narrow and encompasses a significant number of data points.
$\text{MPIW}=\frac{1}{pN}\sum_{j=1}^{p}\sum_{i=1}^{N}(U_{ij}-L_{ij}),$ (18)
$\text{PICP}=\frac{1}{pN}\sum_{j=1}^{p}\sum_{i=1}^{N}\mathcal{I}(U_{ij}<\hat{y}_{ij}<L_{ij}),$
$L_{ij}$ and $U_{ij}$ correspond to the lower and upper bound of the confidence interval for observation road $i$ at time slot $j^{th}$. The $\mathcal{I}(U_{ij}<\hat{y}_{ij}<L_{ij})$ equals 1 if the condition is true, else 0.
4.3.3 Zero Accident Identification Metrics
Given the problem of zero-value data sparsity, highlighted in Figure 2, it is essential to assess the ability of a model to accurately represent roads without accidents. We used the true zero rate (ZR) as a metric to quantify whether a model reflects the zero sparsity seen in the actual data. The formula is as follows:
$\text{ZR}={\frac{1}{pN}\sum_{j=1}^{p}\sum_{i=1}^{N}\mathcal{I}(y_{ij}=0\cap\hat{y}_{ij}=0)},$ (19)
where the indicator function, $\mathcal{I}(y{ij}=0\cap\hat{y}{ij}=0)$, equals 1 if the condition holds true, and 0 otherwise.
4.3.4 Accuracy Hit Metric
To evaluate the effectiveness of our model in identifying roads with a higher likelihood of accident occurrence, we use the Accuracy at $a$ (Acc@$a$) metric. Specifically, we focus on the top 20% ($a=20\%$) of roads considered high-risk to verify if these areas indeed correspond to actual accidents. This alignment of predicted high-risk areas with real accident sites, as reported in Zhang and Cheng, (2020); Zhang et al., 2019b , allows Acc@$a$ to effectively measure the precision of our model in forecasting roads at increased risk of accidents. The formula for calculating Acc@$a$ is as follows:
$Acc@a=\frac{1}{p}\sum_{j=1}^{p}\frac{\sum_{i=1}^{N}\mathcal{I}(y_{ij}>0\cap\hat{y}_{ij}\geq\text{Top}(\hat{y}_{:j},a))}{\sum_{i=1}^{N}\mathcal{I}(y_{ij}>0)},$ (20)
The indicator function, $\mathcal{I}(y_{ij}>0\cap\hat{y}{ij}\geq\text{Top}(\hat{y}{:j},a))$, is rated as 1 when an accident (denoted by $y_{ij}>0$ for an occurrence on the $i^{th}$ road at the $j^{th}$ time slot) aligns with the top $a$ percent of predicted risks (where $\hat{y}{ij}$ meets or exceeds the threshold $\text{Top}(\hat{y}{:j},a)$), and 0 otherwise.
4.4 Baseline Models
In evaluating our proposed model, we carefully select baseline models that allow evaluation of the features of the proposed model: uncertainty quantification in multistep road-level accident risk prediction. Our selection criteria favour traditional point estimation methods and the most recent interval prediction models. This approach is in line with recent studies Zhuang et al., (2022). Specifically, our baselines include the mean-predict method (HA) and advanced spatial-temporal graph deep learning techniques such as STGCN and STGAT. Additionally, we consider parameterisation approaches such as STG-GNN, STNB-GNN, STTD-GNN, and STZINB-GNN. These probabilistic models have two or three parameter distributions, simpler than the four-parameter models we propose. The comparison models are:
  • HA(Liu and Guan, (2004)): The historical average serves as the minimum baseline. It is calculated by averaging road traffic accidents in the same time intervals from historical data to predict the future road traffic accident.
  • STGCN(Yu et al., (2018)): Spatio-temporal graph convolution network, which utilises graph convolution and 1D convolution to capture spatial dependencies and temporal correlations, respectively.
  • STGAT(Veličković et al., (2017)): The Graph Attention Network (GAT) can focus on the features of neighbouring nodes by assigning different attention weights to each node within a neighbourhood. We replaced the GCN with GAT in the STGCN and named the modified model STGAT.
  • STG-GNN(Wang et al., (2023)): Same as STZITD-GNN, we replace the statistical assumption of road traffic accidents with Gaussian distribution.
  • STNB-GNN(Zhuang et al., (2022)): Same as STZITD-GNN, we replace the statistical assumption of road traffic accidents with Negative Binomial Distribution.
  • STTD-GNN: Same as STZITD-GNN, we replace the statistical assumption of road traffic risks with TD Distribution.
  • STZINB-GNN(Zhuang et al., (2022)): Unlike STZITD-GNN, we replace the statistical assumption of road traffic accidents with a zero-inflated negative binomial distribution.
4.5 Performance Comparison
We evaluated the performance of our proposed model, STZITD-GNN, against several baseline models from four different aspects: accuracy of point estimation, uncertainty quantification, non-risk road identification and accurate prediction of accident occurrence, as shown in Tables 4. The results are indicated by upward arrows, representing improvements over the second-best models. These enhancements confirm the significant impact of probabilistic assumptions on the accuracy and stability of our results, consistent with recent findings by Wang et al., (2023). The consistent performance of the STZITD-GNN model underscores the critical importance of selecting appropriate distributional assumptions that align with the characteristics of traffic accident data.
For point estimation accuracy, the superior performance of the STZITD-GNN model across all regression error metrics highlights the effectiveness of the ZITD distribution. This distribution captures the long-tail dynamics of road traffic accident risk data with greater accuracy, making the mean value derived from the four parameters of the distribution decoder more representative of the actual risk scores. STZITD-GNN significantly improves the accuracy of the estimation of the accident risk score, demonstrating a 49. 06% increase in MAPE compared to the second-best model in the Lambeth case study, which has the most significant zero-inflation. Traditional deterministic models like HA, STGCN, and STGAT show varied performance, with HA consistently underperforming. GAT-based models, utilising predefined adjacency matrices that embed domain spatial knowledge, generally outperform GCN models by incorporating spatial proximity and contextual similarity, thereby better representing spatial dependencies. Among probabilistic models, those that assume Gaussian distributions perform poorly. They are significantly outperformed by deterministic models and show far worse performance 50% compared to the STZITD-GNN model, which uses skewed distribution assumptions. This underscores the importance of selecting appropriate distributional assumptions for probabilistic models.
Second, the variation in uncertainty quantification across different models highlights the significant impact of distributional assumptions. The four-parameter distribution model (STZITD-GNN) shows substantial improvements, most notably in effectively capturing a narrow distribution while encompassing a wider range of actual accident risk values. In contrast, the worst-performing model, STG-GNNs, illustrates how improper distributional assumptions can degrade the quality of the estimation, introducing greater uncertainty into the results. The MPIW of this model is approximately 1.5 times that of the least uncertain three-parameter model, due to its constraints that enforce equal mean and variance in the distribution. Although ideally, outcome distributions would have a narrow interval, traditionally a wider MPIW suggests a higher PICP, indicating a trade-off in measuring uncertainty(Wang et al.,, 2023; Zhu et al.,, 2022). The STZINB-GNN and STTD-GNN models show that in scenarios characterised by extreme zero-inflation and heavier tails, a higher MPIW is associated with better PICP performance. This correlation suggests that broader prediction intervals, which encompass more observations, are beneficial in more extreme cases as they cover a wider range of observed values, despite reducing the ability to achieve narrower intervals. However, the STZITD-GNN model shows a promising improvement in MPIW, ranging from 47.30% to 55.07% in all case studies, highlighting its efficacy. In terms of PICP, it also shows a marginal increase of approximately 0.2% over the next best model, STTD-GNNs. This modest improvement still validates the impact of leveraging long-tailed and zero-inflated TD distributions, which contribute to a larger PICP and reduced variance in the predicted outcomes.
In scenarios characterised by a high prevalence of non-accident roads, the STZITD-GNN model not only aims to precisely quantify accident risk values but also effectively differentiates non-accident roads. The integration of the sparsity parameter ($\pi$) significantly enhances the ability of the traditional TD assumption to address the prevalent zeros in the dataset, thereby reducing uncertainties associated with zero outcomes. The TD and ZINB distributions provide distinct parametric approaches to tackle the issue of zero-inflation. The TD distribution naturally integrates zero-inflation through its end-to-end structure, while the ZI parameter in the ZINB model requires an additional interpretative step like ZITD model does. For example, in the Tower Hamlets and Westminster cases, the ZI parameter significantly increases performance in zero-rate predictions compared to the NB distribution, showing improvements of 16.54% and 17.94%, respectively. In extreme zero-inflation scenarios, such as in Lambeth, both the STTD and STZINB models perform poorly, with a 15.09% discrepancy in ZR results compared to the STZITD-GNN model. Although the ZI paradigm is integrated into both the TD and the NB models, the ZITD model has a continuous parameter space, which can be advantageous for fitting data with a mix of zero inflation and continuous positive values. In contrast, traditional ZINB models operate under a discrete probabilistic assumption and struggle with complex data structures, particularly when the positive values are continuous rather than merely counts.
In analyses that correlate high-risk roads with actual accidents, the STZITD-GNN model consistently demonstrates robust performance. In particular, it achieved a significant 76.59% precision in identifying actual accidents in the Lambeth case. Across all three distinct cases, the model maintains a robust and accurate identification rate, with a variance of only 7.61%. In contrast, models employing unsuitable distribution assumptions, such as STG-GNN, tend to significantly underpredict accident occurrences, often yielding the poorest results. This discrepancy underscores the importance of selecting the appropriate distribution model again.
Table 4: Comparison of multi-step prediction performance (14 days) in three boroughs. The bold font indicates the best performing model, while underline indicates the second best. To maintain clarity, detailed names of all GNNs in the final five statistically integrated deep learning models are omitted. Instead, abbreviations such as STG, STNB, STTD, STZINB, and STZITD are used to emphasise the most significant modifications within the GNN models.
Metric Model HA STGCN STGAT STG STNB STTD STZINB STZITD
Dataset Lambeth
Point Estimation MAE 0.1348 0.0613 0.0302 0.1182 0.0802 0.0315 0.0544 $\mathbf{0.0238}_{\uparrow 21.19\%}$
MAPE 0.4142 0.4023 0.3203 0.4046 0.0363 0.0426 0.0265 $\mathbf{0.0135}_{\uparrow 49.06\%}$
RMSE 0.2107 0.1436 0.1122 0.1827 0.1390 0.1098 0.1191 $\mathbf{0.0947}_{\uparrow 13.75\%}$
Uncertainty Quantification MPIW / / / 0.4871 0.1946 0.1349 0.0454 $\mathbf{0.0204}_{\uparrow 55.07\%}$
PICP / / / 0.7328 0.8903 0.9873 0.9796 $\mathbf{0.9899}_{\uparrow 00.26\%}$
Non Accident Quantification ZR 0.5203 0.5305 0.5986 0.1992 0.5621 0.6838 0.6408 $\mathbf{0.7870}_{\uparrow 15.09\%}$
Accurate $Acc@20$ 0.4520 0.6113 0.6422 0.2666 0.4471 0.7123 0.6184 $\mathbf{0.7659}_{\uparrow 07.52\%}$
Dataset Tower Hamlets
Point Estimation MAE 0.0939 0.2385 0.0305 0.1741 0.2475 0.0433 0.2135 $\mathbf{0.0271}_{\uparrow 11.15\%}$
MAPE 0.2896 0.2017 0.1205 0.2662 0.0399 0.0246 0.0594 $\mathbf{0.0221}_{\uparrow 10.16\%}$
RMSE 0.1987 0.1325 0.1355 0.3896 0.1548 0.1186 0.1368 $\mathbf{0.0918}_{\uparrow 22.60\%}$
Uncertainty Quantification MPIW / / / 0.5321 0.2263 0.0810 0.0296 $\mathbf{0.0156}_{\uparrow 47.30\%}$
PICP / / / 0.6439 0.9442 0.9845 0.9845 $\mathbf{0.9871}_{\uparrow 00.26\%}$
Non Accident Quantification ZR 0.4998 0.4896 0.5234 0.1351 0.5158 0.6420 0.7526 $\mathbf{0.8876}_{\uparrow 17.94\%}$
Accurate $Acc@20$ $0.4752$ 0.5869 0.6950 0.2647 0.5022 0.6368 0.5827 $\mathbf{0.7224}_{\uparrow 03.94\%}$
Dataset Westminster
Point Estimation MAE 0.1467 0.0784 0.0847 0.1055 0.3266 0.0579 0.3240 $\mathbf{0.0357}_{\uparrow 38.34\%}$
MAPE 0.1884 0.8976 0.6012 2.9295 0.0330 0.0973 0.0415 $\mathbf{0.0259}_{\uparrow 21.52\%}$
RMSE 0.5999 0.2384 0.2433 0.6681 0.2909 0.1552 0.2751 $\mathbf{0.1015}_{\uparrow 34.60\%}$
Uncertainty Quantification MPIW / / / 0.8542 0.6272 0.2021 0.0493 $\mathbf{0.0259}_{\uparrow 47.46\%}$
PICP / / / 0.6036 0.9363 0.9801 0.9796 $\mathbf{0.9893}_{\uparrow 00.94\%}$
Non Accident Quantification ZR 0.5981 0.1431 0.2922 0.2071 0.5393 0.5079 0.6288 $\mathbf{0.7328}_{\uparrow 16.54\%}$
Accurate $Acc@20$ 0.4217 0.5020 0.4808 0.2933 0.4503 0.6075 0.5139 $\mathbf{0.6898}_{\uparrow 13.55\%}$
4.5.1 Visualization of Predicted Road-Level Accident
To illustrate the results, we have generated choropleth maps of road-level accidents across three distinct boroughs over a 14-day period. Considering the sparsity of accident data, which complicates the visualisation, we selected the three most risk-intensive periods of the data for display, as shown in Figure 4. Each borough reveals distinct accident patterns that correlate closely with the model’s predictions, offering a granular view of its predictive accuracy.
These visualizations highlight the general sparsity of traffic accident data, with most roads being non-risk (blue). Only a few roads exhibit higher accident risk levels (red), while others are associated with low-risk accidents. The STZITD-GNN model excels in identifying a range of risk levels, from zero-risk to high-risk. It performs particularly well in Westminster, accurately identifying a broader range of risk-prone roads, aided by the area’s lower zero inflation rate and denser graph structure. Furthermore, the model consistently and accurately maps the spatial connectivity of high-risk accident roads across all three boroughs, with risk areas typically clustering along major thoroughfares such as A11 Whitechapel Road and A13 in Tower Hamlets and A3 in Lambeth.
In the three distinct boroughs, the STZITD-GNN model’s ability to handle diverse spatial patterns and variations in daily accidents demonstrates its adaptability and consistent performance. However, the model occasionally misclassifies zero-risk roads as low-risk, particularly in Lambeth and Tower Hamlets. Although these errors generally fall within lower-risk categories and do not significantly impact the identification of higher-risk roads with actual accident occurrences, refining this aspect of the model could further improve its overall accuracy.
Figure 4: Predicted long-term (14-day) accidents on Lambeth, Tower Hamlets, and Westminster.
4.6 Sensitivity Analysis
4.6.1 Temporal Stability of Metrics
Evaluating temporal stability is crucial for evaluating a model’s performance over multiple prediction time steps, particularly in data-driven methods, where superior results in specific instances might mask potential underfitting issues (Ding et al.,, 2022; Mannering,, 2018). Our primary goal is to achieve consistently low uncertainty, as indicated by the MPIW metric in three boroughs, depicted in Figure 5. We strive to maintain a narrow yet comprehensive MPIW throughout the prediction steps to ensure reliable prediction intervals for traffic accidents.
Analysis shows that the MPIW of the top performing models, including STZITD-GNNs, is significantly lower than that of the baseline models, emphasising their superior performance throughout the study period. Therefore, we focus on the temporal dynamics of these leading models, presenting baseline performances as averages for contextual clarity.
Figure 5: Temporal variation of MPIW for baseline and STZITD-GNN models.
As indicated in Figure 5, the MPIW is typically lowest on the first day of testing and generally increases over time. However, this increase is smoother in our STZITD-GNNs model compared to the more erratic fluctuations observed in baseline models. This smoother increase reflects our model’s ability to handle the inherent uncertainties and the variability of risk levels, which might be less accurately captured by models with misspecified distributions. Moreover, boroughs like Westminster and Lambeth, known for higher instances of zero-accident risk values, show more fluctuating MPIWs, suggesting greater uncertainties in prediction intervals due to the sparsity of traffic accidents. Despite these challenges, our model consistently delivers stable and accurate predictions across different boroughs.
In the evaluation of the temporal stability for the estimation of points, depicted in Figure 6, the STZITD-GNN customised approach emphasises the advantages of applying the most appropriate probabilistic assumptions, enhancing spatiotemporal representations, and managing the various aspects of traffic accidents. It is apparent that the STG-GNN model consistently shows the weakest performance across all temporal prediction periods, with deterministic models also performing poorly. Furthermore, models based on the Negative Binomial (NB) distribution exhibit unsatisfactory performance characterised by higher fluctuation and higher error rates. This inadequacy stems from the focus of NB models on binary accident-occurrence outcomes, which often overlook the complexities of long-tail accident risk values. In contrast, our model exhibits significantly less variance across all three evaluated metrics at each time step, demonstrating stability even toward the end of the prediction horizon. This robust performance is attributed to the model’s ability to dynamically fine-tune prediction intervals using four critical parameters, effectively addressing the challenges posed by the sparsity and skewed distribution of traffic accident data.
(a) MAE over Time Performance
(b) MAPE over Time Performance
(c) RMSE over Time Performance
Figure 6: Time-wise Comparison for the Point Estimation Metrics
The final component of the evaluation focusses on quantifying both non-accident roads and high-risk roads where real accidents occurred, as shown in Figure 7. It is evident that the STZITD-GNN model consistently excels in both metrics, maintaining stable performance with less than 18% variance throughout the 14-step period. As illustrated in Figure 7(a), deterministic models exhibit less variability but face substantial challenges in accurately identifying true nonaccident roads, often misclassifying all roads as having no accident risk in extreme zero inflation scenarios, leading to overfitting. Conversely, probabilistic models employing Gaussian distributions struggle to accommodate the high prevalence of zero values, resulting in poor performance and high sensitivity in the ZR metrics over time. This issue arises because each time step may present unique scenarios of zero-valued roads, which Gaussian models often misinterpret due to their uniform mean and variance assumptions.
In the identification of High-Risk Accident Roads as illustrated in Figure 7(b), the ZITD distribution particularly excels. It incorporates a weighted dispersion parameter ($\phi$) to distinguish the lower accident risk from the non-accident road, displaying the best of low, middle and high risk roads with long-tailed distributions. This adaptability to the complex temporal dynamics of road-level accident data at each time step not only mitigates overfitting, but also stabilises the model performance, enabling it to accurately capture real accident occurrences in critical geographic locations.
(a) Zero Rate over Time Performance
(b) Accuate Hit at Top 20% Risky Roads over Time Performance
Figure 7: Time-wise Comparison for the Non Accident Roads Quantification and Accurate Accident Hit Metrics
4.6.2 Analysis of Long-tail Accident Data Characteristics
To explore the impact of inherently long-tailed data on our model, we created a three-dimensional surface plot over four information sources, as illustrated in Figure 8. Due to visualisation constraints, it is not feasible to represent all four parameters in a single plot. We therefore focus on the TD parameters $\phi$, $\rho$, and $\mu$, with the understanding that the sparsity parameter $\pi$ significantly influences the original TD parameters. This plot illustrates the predicted accident values, with the actual accident values $y$ represented by the colour gradient.
A distinct trend is evident in the learnt values of $\rho$, consistently approaching 2, suggesting a preference for more skewed distributions as $\rho$ increases from 1 to 2 (Tweedie et al.,, 1984; Kurz,, 2017). This trend corresponds with the long-tailed nature of our data and validates the probabilistic assumptions. The increase in $\rho$ serves as an adaptive response to the component of the loss function $y\frac{\mu^{1-\rho}}{1-\rho}(y>0)$, effectively penalising the underestimation of true, although infrequent, higher values of accident risk within the long tail. This adjustment directs the model to focus more on such kinds of traffic accident rates, continually benefiting from significant enhancements in model sensitivity and accuracy.
Additionally, significant variations in the values $\phi$ are observed, particularly concerning data with zero value, which serve as the dispersion parameter. For example, in Westminster, $\phi$ is significantly lower than in other boroughs, indicating fewer zero-inflation issues. Traditionally, $\phi$ in the TD distribution is crucial for managing the frequency of zero occurrences to effectively address zero inflation. However, due to the highly imbalanced nature of our data, a single $\phi$ parameter often fails to capture the prevalence of zeros, as demonstrated by the overlap of plots in areas of lower and zero accident values. In our ZITD model, an enhanced $\phi$, augmented by the sparsity parameter $\pi$, manages these features, offering a precise depiction of the inherent sparsity and discrete uncertainty in road-level accident predictions.
Moreover, the model parameter $\mu$, which denotes the mean of the predicted accident values, shows a strong correlation with the actual accident values $y$, as evidenced by the colour alignment in the graphs. This correlation not only substantiates the effectiveness of the model in estimating mean points, but also affirms its capability to accurately mirror and adapt to the complexities of urban accident data.
(c) Westminster
(d) Lambeth
(e) Tower Hamlets
Figure 8: Three Tweedie parameter visualization on Lambeth, Tower Hamlet, and Westminster with multi-step predict. The colour bar shows the predicted accident from the highest to zero.
5 Conclusion and Discussion
This paper introduces STZITD-GNN, a probabilistic deep learning framework that, to the best of our knowledge, is the first uncertainty-aware model specifically tailored for road-level accident prediction. Developed from data-driven insights, the STZITD-GNN utilises predefined distribution characteristics of traffic accident data to enhance the representation of sparse traffic accident scenarios. It effectively detects finer-level correlations in traffic accidents, thereby delivering precise predictions with narrow prediction intervals and reduced uncertainties. Through systematic multi-timescale experiments conducted on three real-life urban accident datasets, our results demonstrate that the efficacy of our STZITD-GNN approach surpasses that of state-of-the-art probabilistic deep learning models as well as deterministic models. The key contributions of this study are summarised as follows.
  • Innovative Distribution Extension: We have adapted the conventional TD to a ZITD distribution, introducing a sparsity parameter, $\pi$, integrated with the dispersion parameter $\phi$. This modification tailors the model to our data-driven assumptions, representing the dispersed and extreme zero-inflated nature of accident data at the road level.
  • Integration with ST-GNN: The incorporation of the ZITD encoder into the ST-GNN framework makes the STZITD-GNN model an end-to-end framework to handle the complex spatiotemporal dynamics of accidents. It surpasses traditional point estimation methods by enabling the mapping of output variance and prediction intervals. Additionally, the model facilitates an interpretable, uncertainty-aware process that elucidates the underlying factors contributing to improved prediction accuracy.
  • Empirical Validation Across Boroughs:The model has shown improved predictive accuracy and minimal uncertainties compared with benchmark models in three London boroughs, each featuring unique spatial accident patterns in multi-step scenarios. This performance confirms the robustness and generalizability of the model across diverse urban settings, even with sparse datasets. Additionally, the model’s stable performance across heterogeneous datasets, which vary in accident determination criteria, further underscores its reliability and effectiveness in urban traffic risk assessment.
It should be recognised that the challenges of overdispersion and zero inflation extend beyond traffic accidents, affecting a wide array of urban issues such as crime rates. This universality suggests that our proposed model could be applicable to a broader range of contexts beyond its current usage.
While our research demonstrates its effectiveness in predicting traffic accidents, there are several limitations that provide direction for future research. We aim to enhance the accuracy of mapping accidents at finer temporal resolutions and plan to expand the model’s applicability to various countries under consistent assumptions. Future developments will focus on creating a more adaptable model that can accommodate geographic differences, thus increasing its usability across diverse regions. In addition, we intend to integrate more diverse data types into our model. Incorporating GPS data, for example, could provide more detailed insight into traffic exposure. To further improve the accuracy and utility of our model, we also plan to use more comprehensive data sets, including urban characteristics derived from street or satellite images. This would enrich our model’s contextual understanding and could identify factors that exacerbate traffic accident impacts across different locations and times. Addressing these challenges in future studies will improve the robustness and applicability of the framework in real-world scenarios, ultimately contributing to the development of more effective traffic management systems and accident prevention strategies.
In conclusion, the development of the STZITD-GNN model marks a significant advancement in the prediction of urban road-level traffic accidents. However, this is just the beginning of an ongoing effort to refine and enhance this model. By integrating additional data sources and continually improving our methodologies, our goal is to develop a tool that not only mitigates traffic risks but also significantly contributes to global urban safety strategies.
6 Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work the author(s) utilized GPT-4 for preliminary grammar checks and proofreading. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.
Appendix A Tweedie Table
Tweedie EDMs $\rho$ $\kappa(\theta)$ $\phi$ $\theta$ $a(y,\phi)$
Normal 0 $\frac{\theta^{2}}{2}$ $\sigma^{2}$ $\mu$ $\exp{(-y^{2})/2\sigma^{2}-\log 2\pi/2)}$
Poisson 1 $\exp{(\theta)}$ $1$ $\log\mu$ $1/y!$
Poisson-Gamma (1,2) $\frac{(\theta-\rho\theta)^{\frac{2-\rho}{1-\rho}}}{2-\rho}$ $\phi$ $\frac{\mu^{1-\rho}}{1-\rho}$ Eq.5
Gamma 2 $-\log(-\theta)$ $\phi$ $-1/\mu$ $\frac{\phi^{-1/\phi}y^{1/\phi-1}}{\Gamma(1/\phi)}$
where $\mu,\sigma$ is the mean and the standard values of Gaussian Distribution.
Appendix B Temporal encoder – GRU
Traffic accidents demonstrate significant temporal correlation, underscoring the idea that the dynamic temporal features of roads encapsulate a wealth of information. This observation has been acknowledged and exploited in prior methodologies (Wang et al., 2021a, ; Zhao et al.,, 2019). As shown in Figure 1, to seize these temporal proximity features, we implement a GRU (Chung et al.,, 2014). The GRU, in comparison to traditional LSTM networks, offers the advantage of learning relatively long-term dependencies without facing issues of vanishing and exploding gradients. In the GRU framework, there isn’t a segregation of internal and external states, as seen in LSTM networks. Instead, it addresses the gradient vanishing and exploding problem by directly integrating a linear dependency between the current network state $h_{t}$ and the preceding state $h_{t-1}$. Consequently, while a GRU preserves the capabilities of an LSTM network, its architecture remains more streamlined (Gu et al., (2019); Mahmoud et al., (2021); Ma et al., (2022)).
Specifically, this component accommodates the feature sequence over past time windows $1:t$, thereby learning temporal dependencies. We represent the temporal embedding as $\mathcal{Z}_{T}$, taking into account both historical spatiotemporal road features and traffic accidents as input:
$\mathcal{Z}_{T}=\text{GRU}(X_{1:t},Y_{1:t}),$ (21)
Once the input features have been processed via the GRU, we obtain historical temporal embeddings $\mathcal{Z}T$ for all roads. These are subsequently directed to graph neural encoders. The GRU, in particular, consists of two gates: the update gate and the reset gate. The encoder processes a single input $X_{t},Y_{t}$ at time slot $t$ and prior hidden features $h_{t-1}$ in sequence:
$r_{t}=\sigma(W_{r}\cdot[h_{t-1},X_{t},Y_{t}]+b_{r}),$ (22)
$u_{t}=\sigma(W_{u}\cdot[h_{t-1},X_{t},Y_{t}]+b_{u}),$
$\tilde{h}_{t}=\text{tanh}(W_{c}\cdot[r_{t}\cdot h_{t-1},X_{t},Y_{t}]+b_{c}),$
$h_{t}=(1-u_{t})\cdot h_{t-1}+u_{t}\cdot\tilde{h}_{t}$
In the equation above, $h_{t}$ and $h_{t-1}$ represent hidden features at the current and preceding time steps, respectively, while $\tilde{h}_{t}$ denotes the candidate state that temporarily stores information from reset and update gates ($r_{t}$,$u_{t}$). The variables $W$ and $b$ (i.e. $W_{r},W_{u},W_{c},b_{r}$, $b_{u},b_{c}$) represent learnable weight and bias matrices, respectively. Importantly, $r_{t}$ acts as the reset gate, determining the extent of integration of hidden features from previous time steps into features at the current time step. Meanwhile, $u_{t}$ serves as the update gate, controlling how much information from previous time steps will be carried over to the next time step. The $\text{tanh}(\cdot)$ is the hyperbolic tangent function. We perform the GRU gate operations step by step and select the final layer’s hidden features $h_{t}$ as the temporal embedding $\mathcal{Z}_{T}\in\mathbb{R}^{N\times F}$, where $F$ denotes the dimension of road temporal embedding.
Appendix C Spatial encoder – GAT
Within our research, road data is fundamentally perceived as having an inherent graph structure. This perspective facilitates the interpretation of node features as graph signals. To accommodate such graph-structured information, we employ the GAT (Veličković et al.,, 2017), an effective methodology renowned for enabling potential model deepening while concurrently incorporating dynamic spatial correlation into our network, achieved through real-time adjustment of attention weights in response to data changes (Wang et al.,, 2022; Zhang et al., 2019a, ).
Distinct from GCN of a spectral-domain convolution (Kipf and Welling, 2016a, ), GAT operates as a spatial-domain convolutional network. Instead, GAT emphasizes the magnitude of the graph signal value at distinct points and their inter-distance to execute the convolution process. This trait empowers GAT to alleviate the over-smoothing issue associated with GCN deepening, thereby augmenting prediction precision regarding peaks and valleys (Zhao et al.,, 2019).
Initially, we compute the attention coefficient $\alpha_{ij}$ between roads $v_{i}$ and $v_{j}$ as per the following equation:
$\alpha_{ij}=\frac{\exp({\text{LeakyReLU}\bigl{(}\vec{a}^{\mathsf{T}}[W_{a}\mathcal{Z}{T_{i}}||W_{a}{\mathcal{Z}{T_{j}}])\bigl{)}}}}{\sum_{A_{i,o}>0}\exp({\text{LeakyReLU}\bigl{(}\vec{a}^{\mathsf{T}}[W_{a}{\mathcal{Z}{T_{i}}}||W_{a}{\mathcal{Z}{T_{o}}}])\bigl{)}}},$ (23)
In this equation, $A_{i,o}>0$ represents the neighbourhood or connected roads of the road $v_{i}$. $W_{a}\in\mathbb{R}^{F\times F^{\prime}}$ and $\vec{a}\in\mathbb{R}^{2F^{\prime}}$ are learnable weights, where $F^{\prime}$ denotes the hidden dimension. We apply the concatenation function [||][$\cdot$|$\cdot$][ ⋅ | | ⋅ ] to concatenate the embedding of $v_{i}$ and its neighbours, utilizing learnable parameter $W_{a}$ and mapping function $\vec{a}$. Subsequently, we employ an activation function, $\text{LeakyReLU}(\cdot)$, and normalize the attention coefficient $\alpha_{ij}$ using the $\exp(\cdot)$ function. The embedding of road $v_{i}$ is consequently updated with a linear combination of the neighbourhood of road $v_{i}$ and itself, wherein these computed attention coefficients are employed.
In order to preemptively address potential inaccuracies in attention assignments, we have incorporated multiple attention assignments, otherwise known as multiple attention heads, into our approach. This tactic greatly amplifies the expressive capacity of the model. The multi-head attention mechanism facilitates concurrent consideration of information from a wide range of representation subspaces at various positions. This enables each head to learn and focus on different facets of the input data, thereby enhancing the model’s ability to comprehend and accurately represent the data (Reza et al.,, 2022; Tang and Zeng,, 2022; Ding et al.,, 2023). Herein, $M$ symbolizes the number of attention mechanisms, with the output embeddings from these mechanisms being concatenated in the following manner:
$\mathcal{Z}_{i}=||_{m=1}^{M}\sigma\left(\sum_{A_{i,j}>0}\alpha_{ij}^{m}{W_{a}}^{m}\mathcal{Z}_{T_{j}}\right),$ (24)
In this equation, superscript $m$ denotes a different head. The embedding of the road $v_{i}$ is optimized by aggregating its neighbours’ embedding with attention coefficients $\alpha_{ij}^{m}$, and we apply learnable weights ${W_{a}}^{m}$ for different heads. After this, we apply a nonlinear sigmoid function and concatenate different heads of embedding. The multi-head attention mechanisms hereby ensure stability in the attention mechanisms learning process and encapsulate diverse perspectives of road-level graph features. Notably, in Eq 24, the dimension of the output feature is $M\times F^{\prime}$, rather than $F^{\prime}$. To ensure consistency in the output feature’s dimension, we compute the average of each road’s embedding in the last layer as follows:
$\mathcal{Z}_{i}=\sigma\left(\frac{1}{M}\sum_{m=1}^{M}\sum_{A_{i,j}>0}\alpha_{ij}^{m}{W_{a}}^{m}\mathcal{Z}_{T_{j}}\right).$ (25)
Finally, we input temporal embedding $\mathcal{Z}_{T}$ into GAT to capture spatial dependency and obtain spatiotemporal $\mathcal{Z}_{i}\in\mathbb{R}^{F^{\prime}}$ of road $v_{i}$. After that, we use $\mathcal{Z}$ to infer the four-parameter STZITD-GNNs model.